OpenStudy (nbarrera):

Help+Medal The expression (2x + 3)(3x + 4) simplifies to which of the following? A. 2x(3x + 4) + (3x + 4) B. 2x(2x + 3) + 3(3x + 4) C. 2x(3x + 4) + 3(3x + 4) D. 2x(3x + 4) The answer i picked is C but i just want to make sure if i'm right.

OpenStudy (anonymous):

Yes, C is right.

OpenStudy (nbarrera):

really?

OpenStudy (anonymous):

It's the distributive law: \[(a+b)\cdot c = ac + bc\] And in this case: \[a = 2x, b = 3, \text{and} \, c = (3x+4)\]